OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 2x2 - 4x - 1

OpenStudy (anonymous):

or i mean f(x) = 2x^2 - 4x - 1

OpenStudy (whpalmer4):

Find the values of x where f(x) = 0. The axis of symmetry is x = the average of those two. Or, as you have the equation in the form \(y = ax^2+bx+c\), the x value of the vertex (which is on the line of symmetry) can be gotten from \(x = -\frac{b}{2a}\)

OpenStudy (anonymous):

so it would be 1 correct?

OpenStudy (whpalmer4):

\(b = -4,~a = 2,~-\frac{b}{2a} = -\frac{-(-4)}{2*2} = 1\) Yes, 1 is correct.